OpenStudy (anonymous):

im trying to do my math hw but i dont get it we are studying Quadratic functions abd the function is y=x^2-10x+2 help please :((

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

First of all: Don't call yourself a math retard

myininaya (myininaya):

lol

OpenStudy (anonymous):

haha (:

myininaya (myininaya):

the first step is do not be negative lol

OpenStudy (anonymous):

haha

myininaya (myininaya):

what would you like to do with the "quady"

OpenStudy (anonymous):

lol ok ok but can i get some gelp :/

OpenStudy (anonymous):

*help

OpenStudy (anonymous):

Second of all: Its just like solving an algebraic equation

myininaya (myininaya):

i like gelp better

OpenStudy (anonymous):

same lol

OpenStudy (anonymous):

lol i know its the same but i juss dnt get it

myininaya (myininaya):

so do you want to graph the "quady"?

myininaya (myininaya):

silence satellite!

OpenStudy (anonymous):

i dnt think im supposed to graph i think i juss needa find the vertex and coordinates :/ or maybe i am idk >.<

OpenStudy (anonymous):

myininaya izzat you?

myininaya (myininaya):

omg 465 so lame you are leaving me in the dust the more we find the better are graph will look lets start out with x-intercepts

myininaya (myininaya):

to find x-intercepts, set y=0 and solve for x

myininaya (myininaya):

x^2-10x+2=0

OpenStudy (anonymous):

okk then what :o

myininaya (myininaya):

now this is totally not factorble so you have to use the quady formula!

myininaya (myininaya):

\[x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

myininaya (myininaya):

satellite i was kidding you can talk lol

myininaya (myininaya):

\[x=\frac{10 \pm \sqrt{10^2-4(1)(2)}}{2(1)}\]

OpenStudy (anonymous):

sooo after tht wat next o.o

myininaya (myininaya):

\[x=\frac{10 \pm \sqrt{100-8}}{2}=\frac{10 \pm \sqrt{92}}{2}=\frac{10 \pm \sqrt{4*23}}{2}\]

myininaya (myininaya):

\[x=\frac{10 \pm 2\sqrt{23}}{2}=5 \pm \sqrt{23}\]

OpenStudy (anonymous):

lordamercy \[x^2-10x+2=0\] \[x^2-20x=2\] \[(x-5)^2=-2+25=23\] \[x-5=\pm\sqrt{23}\] \[x=5\pm\sqrt{23}\]

myininaya (myininaya):

so these two are where your graph is going to the intercept the x axis

myininaya (myininaya):

satellite used completeing the square (which is used to proved the quady formula is the solution of the quady equation)

OpenStudy (anonymous):

oops typo. second line should have been \[x^2-10x=-2\]

OpenStudy (anonymous):

also it is nice to use when the middle coefficient is even so you do not have to simplify the radical, factor and cancel

myininaya (myininaya):

yes that too

OpenStudy (anonymous):

the quadratic formula forces a denominator on you but you see that there isn't one!

OpenStudy (anonymous):

i will be quiet now

myininaya (myininaya):

now lets take a break now from the hard stuff and see if we can find a y-intercept

myininaya (myininaya):

to find the y-intercept, set x=0 and solve for y

myininaya (myininaya):

y=0^2-10(0)+2=0-0+2=2

OpenStudy (anonymous):

smhh i know you guys are gonna get mad but...i still dnt get it :[

myininaya (myininaya):

i'm furious lol

OpenStudy (anonymous):

i'll just take my frends answers in school tommorow lol

myininaya (myininaya):

i was kidding

myininaya (myininaya):

what part is getting you down?

OpenStudy (anonymous):

i know lols

OpenStudy (anonymous):

the whooooole thing D:

myininaya (myininaya):

do you know the quadratic formula?

OpenStudy (anonymous):

ax^2=bx+c?

OpenStudy (anonymous):

*ax^2+bx+c

myininaya (myininaya):

\[ax^2+bx+c=0\] -> quad equation \[x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\] -> quad formula (solution to the quad equation)

OpenStudy (anonymous):

mhmm :x

OpenStudy (anonymous):

i thought it was just -b/2a

myininaya (myininaya):

no

OpenStudy (anonymous):

wow ok :/

OpenStudy (anonymous):

bt anyway i go sleep now lols so ttyl

myininaya (myininaya):

ok gn

OpenStudy (anonymous):

gn ^_^